OpenStudy (anonymous):

Use parametric differentiation to differentiate the function defined by x = t+1/t and y = t − 1/t, and find the tangent and normal at the point T where t = 2. I found the gradient to be 5/3

OpenStudy (anonymous):

I will tell you how to do parametric differentiation..I am not sure about this second part

OpenStudy (anonymous):

i done the parametric part :P

OpenStudy (anonymous):

i cant find the equation to the tangent :(

OpenStudy (anonymous):

lol..then i am helpless

OpenStudy (anonymous):

@satellite73 @Hero @agent0smith Help please :)

OpenStudy (agent0smith):

For the tangent line you probably want a y=mx+b equation. If you already have the slope (m) then you need to find b, by using the point that the line passes through... use t=2 to find the x and y values, and use those to find b. The normal will be perpendicular to the tangent line, and pass through that same point.

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

we need this equation? y-y1 = m(x-x1)?

OpenStudy (agent0smith):

That works too.

OpenStudy (anonymous):

So how can we find y1 and x1?

OpenStudy (agent0smith):

Err... read the post above :P "use t=2 to find the x and y values, and use those to find b."

OpenStudy (anonymous):

ok thanks :D i got it